OpenStudy (anonymous):

A committee of four is to be selected from 12 people. In how many different ways can this be done if either Mary or stefan must be on this committee but not both. (we have to use nCr notation and we can use cas calc

OpenStudy (cwrw238):

Mary on committee:- = 11C3 Stefan :- 11C3

OpenStudy (cwrw238):

total ways = 2 * 11C3

OpenStudy (anonymous):

yeah then what?

OpenStudy (cwrw238):

well use you calculator to work out 11C3 11C3 = (11*10*9) / (3*2*1)

OpenStudy (cwrw238):

11C3 means the number of combinations of 3 taken from 11

OpenStudy (anonymous):

yes but thats both. the question is asking for one or the other. not both. the answers is 240

OpenStudy (anonymous):

adding those two gives 330 but this include having both mary and stefan doesnt it?

OpenStudy (cwrw238):

Yes - my mistake its 2 * 10C3 = 2 * 120 = 240

OpenStudy (anonymous):

ok seem right. but why are you picking 3 from 10

OpenStudy (anonymous):

oh nvm i get it now. CHeers